OpenStudy (anonymous):

how do i solve this equation 4x-9y=18

OpenStudy (sandra):

hmmm, if you want to solve for both variables, you will need one more equation

OpenStudy (sandra):

if you want to solve with respect to one variable, e.g. put it into slope intercept form, then that can be done

OpenStudy (sandra):

1. 4x - 9y = 18 2. 4x - 9y - 18 = 18 -18 3. 4x -9y + 9y - 18 = 0 + 9y 4. 9y = 4x - 18 5. 9y/9 = 4x/9 - 18/9 6. y = 4/9x - 2